# A Chung-Fuchs type theorem for skew product dynamical systems

### v1

@@ Abstract.
We prove a Chung-Fuchs type theorem for skew product dynamical systems such that for a measurable function on such a system, if its Birkhoff average converges to zero almost surely, and on typical fibres its Birkhoff sums have a non-trivial independent random walk component, then its associated generalised random walk oscillates, that is the supreme of the random walk equals to +$\infty$+ ∞ and the infimum equals to $\infty$- ∞.
@@ 1. Introduction
We have the following well-known theorem of Chung and Fuchs [4]:
@@
If $\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$, then $S$ is recurrent. Furthermore, if $\mathbb{P}(W_{1}\neq 0)>0$, then $S$ is oscillating, that is
(1.1) $\mathbb{P}(\liminf_{n\to\infty}S_{n}=-\infty)=\mathbb{P}(\limsup_{n\to\infty}S_{n}=+\infty)=1.$
In [4], the condition “$\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$” is to ensure that one can apply the strong law of large numbers to obtain
(1.2) $\mathbb{P}(\lim_{n\to\infty}\frac{S_{n}}{n}=0)=1.$
@@
For oscillation (1.1), Chung and Fuchs showed that the set of recurrent points is a closed additive subgroup of $\mathbb{R}$, hence it is either empty, or the whole space $\mathbb{R}$, or a lattice $a\cdot\mathbb{Z}$ for some $a\in\mathbb{R}$. So if $S$ is recurrent, then the set of recurrent points is non-empty, thus it is either $\mathbb{R}$ or $a\cdot\mathbb{Z}$ for some $a\neq 0$ (since $\mathbb{P}(W_{1}\neq 0)>0$), hence either case implies (1.1).
@@
The following extension of Chung-Fuchs theorem to generalised random walks is essentially due to Dekking [5] (see also [2] when assuming $f\in L^{1}(\mu)$). For readers’ convenience we shall present a proof in the next section.
@@
(1.4) $\mathbb{Q}(\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty)=1,$
where, in brief, the probability measure $\mathbb{Q}$ is a coupling of $\mathbb{P}$ and $\mu$, $S_{n}f(\cdot)$ is a generalised random walk under $\mu$ and $S_{n}F(x,\cdot)$ is a random walk under $\mathbb{Q}_{x}$ for $\mu$-a.e. $x$, where $\mathbb{Q}_{x}$ is the disintegration of $\mathbb{Q}$ w.r.t. $\mu$. In the sub-critical case when $S_{n}f+S_{n}F$ has a positive drift, that is $\mathbb{Q}$-a.s. $\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=c>0$, it is easy to deduce (1.4). The critical case when $\mathbb{Q}(\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=0)=1$ corresponds to (1.3) and it becomes more delicate to verify (1.4), which has led us to derive the following result (a slightly more general and more abstract version of the result in [3]).
@@
Consider the skew-product measure-preserving dynamical system
$(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q}),$
where
@@
    $\mathbb{Q}$ is a coupling of $\mu$ and $\mathbb{P}$, that is $\Pi_{X}^{*}(\mathbb{Q})=\mu$ and $\Pi_{\Omega}^{*}(\mathbb{Q})=\mathbb{P}$, where $\Pi_{X}$ and $\Pi_{\Omega}$ are the projection from $X\times\Omega$ onto $X$ and $\Omega$ respectively.
@@
and let $\mathbb{Q}^{\eta}_{x,\omega}$ denote the conditional measure of $\mathbb{Q}$ w.r.t. $\eta$. For $\mathbb{Q}$-a.e. $(x,\omega)$, $\mathbb{Q}^{\eta}_{x,\omega}$ is a probability measure carried by $\{x\}\times\Omega$. Note that $\mathbb{Q}^{\eta}_{x,\omega}$ only depends on $x$. Hence, equivalently, we may define a family of probability measures $\mathbb{Q}_{x}=\Pi_{\Omega}^{*}(\mathbb{Q}^{\eta}_{x,\omega})$ on $\Omega$ for $\Pi_{X}^{*}(\mathbb{Q})=\mu$-a.e. $x\in X$ such that for any measurable function $F:X\times\Omega\to\mathbb{R}$,
@@
    for $\mu$-a.e. $x\in X$, the sequence

### v2

@@ Abstract.
We prove a Chung-Fuchs type theorem for skew product dynamical systems such that for a measurable function on such a system, if its Birkhoff average converges to zero almost surely, and on typical fibres its Birkhoff sums have a non-trivial independent structure, then its associated generalised random walk oscillates, that is the supremum of the random walk equals to +$\infty$+ ∞ and the infimum equals to $\infty$- ∞.
@@ 1. Introduction
Regarding the recurrence of random walks we have the following well-known theorem of Chung and Fuchs [3]:
@@
If $\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$, then $S$ is recurrent. Furthermore, if $\mathbb{P}(W_{1}\neq 0)>0$, then $S$ oscillates, that is
(1.1) $\mathbb{P}\big{(}\liminf_{n\to\infty}S_{n}=-\infty\big{)}=\mathbb{P}\big{(}\limsup_{n\to\infty}S_{n}=+\infty\big{)}=1.$
In [3], the condition “$\mathbb{E}(|W_{1}|)<\infty$ and $\mathbb{E}(W_{1})=0$” is only to ensure that, by the strong law of large numbers,
(1.2) $\mathbb{P}\Big{(}\lim_{n\to\infty}\frac{S_{n}}{n}=0\Big{)}=1.$
@@
For (1.1), Chung and Fuchs showed that the set of recurrent points is a closed additive subgroup of $\mathbb{R}$, hence it is either empty, or the whole space $\mathbb{R}$, or a lattice $a\cdot\mathbb{Z}$ for some $a\in\mathbb{R}$. So if $S$ is recurrent, then the set of recurrent points is non-empty, thus it is either $\mathbb{R}$ or $a\cdot\mathbb{Z}$ for some $a\neq 0$ (since $\mathbb{P}(W_{1}\neq 0)>0$), and either case implies (1.1).
@@
The following extension of Chung-Fuchs theorem to generalised random walks is essentially due to Dekking [4] (see also [1] when assuming $f\in L^{1}(\mu)$).
@@
(1.4) $\mathbb{Q}\Big{(}\limsup_{n\to\infty}S_{n}f+S_{n}F=+\infty\Big{)}=1,$
where, in brief, the probability measure $\mathbb{Q}$ is a skew-product extension of the ergodic measure $\mu$ to $X\times\Omega$, $S_{n}f(\cdot)$ is a generalised random walk under $\mu$ and $S_{n}F(x,\cdot)$ is a random walk under $\mathbb{Q}_{x}$ for $\mu$-a.e. $x$, where the probability measure $\mathbb{Q}_{x}$ on $\Omega$ is the disintegration of $\mathbb{Q}$ w.r.t. $\mu$. In the sub-critical case when $S_{n}f+S_{n}F$ has a positive drift, that is $\mathbb{Q}$-a.s. $\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=c>0$, it is easy to deduce (1.4). The critical case when $\mathbb{Q}(\lim_{n\to\infty}\frac{S_{n}f+S_{n}F}{n}=0)=1$ corresponds to (1.3) and it becomes more delicate to verify (1.4), which has led us to derive the following result (a more general and more abstract version of the result in [2]).
@@
Recall that $(X,\mathcal{B},T,\mu)$ is a m.p.d.s with $\mu(X)=1$. Consider a skew-product measure-preserving dynamical system $(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q})$ where
@@
    $\Pi_{X}^{*}(\mathbb{Q})=\mu$, where $\Pi_{X}$ is the projection from $X\times\Omega$ onto $X$.
@@
and let $\mathbb{Q}^{\eta}_{x,\omega}$ denote the conditional measure of $\mathbb{Q}$ w.r.t. $\eta$. For $\mathbb{Q}$-a.e. $(x,\omega)$, $\mathbb{Q}^{\eta}_{x,\omega}$ is a probability measure carried by $\{x\}\times\Omega$. Note that $\mathbb{Q}^{\eta}_{x,\omega}$ only depends on $x$. Hence, denoting $\Pi_{\Omega}$ the projection from $X\times\Omega$ onto $\Omega$, we may define a family of probability measures $\mathbb{Q}_{x}=\Pi_{\Omega}^{*}(\mathbb{Q}^{\eta}_{x,\omega})$ on $\Omega$ for $\mu$-a.e. $x\in X$ such that for any measurable function $F:X\times\Omega\to\mathbb{R}$,
@@
The family of probability measures $\mathbb{Q}_{x}$ for $\mu$-a.e. $x\in X$ is also referred as the disintegration of $\mathbb{Q}$ with respect to $\mu$.
@@
    for $\mu$-a.e. $x\in X$,
    (i) $F(x,\cdot)$ is not $\mathbb{Q}_{x}$-a.s. a constant;
    (ii) the sequence $\{F(x,\cdot),F\circ\hat{T}(x,\cdot),F\circ\hat{T}^{2}(x,\cdot),\cdots\}$ is an independent sequence of random variables under $\mathbb{Q}_{x}$.
For $n\geq 1$ denote by
$\hat{S}_{n}F=F+F\circ\hat{T}+\cdots+F\circ\hat{T}^{n-1}.$
We have the following extension of Chung and Fuchs theorem.
Theorem 1.3.
Assume (A). If
$\mathbb{Q}\Big{(}\lim_{n\to\infty}\frac{\hat{S}_{n}F}{n}=0\Big{)}=1,$
then
$\mathbb{Q}\Big{(}\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\Big{)}>0\text{ and }\mathbb{Q}\Big{(}\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\Big{)}>0.$
Remark 1.4.
If we further assume that $\mathbb{Q}$ is $\hat{T}$-ergodic, then, since $\{\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\}$ and $\{\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\}$ are $\hat{T}$-invariant sets, Theorem 1.3 implies
$\mathbb{Q}\Big{(}\liminf_{n\to\infty}\hat{S}_{n}F=-\infty\Big{)}=\mathbb{Q}\Big{(}\limsup_{n\to\infty}\hat{S}_{n}F=+\infty\Big{)}=1.$
Remark 1.5.
Assumption (A) holds for a more general class of skew-product measures $\mathbb{Q}$ than the Peyrière measure studied in [2]. In particular we are not requiring that the law of $F(x,\cdot)$ under $\mathbb{Q}_{x}$ to be independent of $x$.
Remark 1.6.
We improved the proofs comparing to [2]. Firstly we used a result of Kesten to remove the method of using the stopping time that $F\circ\hat{T}^{n}(x,\cdot)$ firstly becomes positive, then we added an induced system argument to deduce that a measurable (not necessarily integrable) coboundary $u-u\circ T=v$ with $u,v\geq 0$ is equal to zero almost everywhere.
Example
We may consider the following canonical example: let $(X\times Y,\mathcal{B}_{X\times Y},\tilde{T},\Theta)$ be a skew product dynamical system of $(X,\mathcal{B},T,\mu)$, where the skew product $\tilde{T}(x,y)=(T(x),\varphi_{x}(y))$ is defined for a family of measurable functions $\varphi_{x}:Y\to Y$ and the $\tilde{T}$-invariant measure $\Theta$ satisfies $\Pi_{X}^{*}(\Theta)=\mu$. For $\mu$-a.e. $x\in X$ let $\Theta_{x}$ denote the disintegration of $\Theta$ w.r.t. $\mu$. We define the skew-product dynamical system $(X\times\Omega,\hat{\mathcal{B}},\hat{T},\mathbb{Q})$ as follows: Let $\Omega$ be the infinite product space of $Y$, that is,
$\Omega=\bigotimes_{\mathbb{N}}Y.$
The family of measurable functions $g_{x}:\Omega\to\Omega$ used to define the skew product $\hat{T}(x,\omega)=(T(x),g_{x}(\omega))$ is taken to be the left-shift operator $\sigma$, that is,
$g_{x}(\omega)=\sigma(\omega)=(y_{2},y_{3},\cdots)\text{ for }\omega=(y_{n})_{n\geq 1}\in\Omega.$
The $\hat{T}$-invariant measure $\mathbb{Q}$ on $X\times\Omega$ is defined as
$\mathbb{Q}(\mathrm{d}x,\mathrm{d}\omega)=\mu(\mathrm{d}x)\mathbb{Q}_{x}(\mathrm{d}\omega),$
where the fibre measures $\mathbb{Q}_{x}$ for $\mu$-a.e. $x\in X$ is the product measure
$\mathbb{Q}_{x}(\mathrm{d}\omega)=\prod_{k=1}^{\infty}\Theta_{T^{k-1}(x)}(\mathrm{d}y_{k})\text{ for }\omega=(y_{n})_{n\in\mathbb{N}}\in\Omega.$
Let $h:X\times Y\mapsto\mathbb{R}$ be a measurable function. Define the measurable function $F:X\times\Omega\to\mathbb{R}$ by
$F(x,\omega)=h(x,y_{1})\text{ for }x\in X\text{ and }\omega=(y_{n})_{n\in\mathbb{N}}\in\Omega.$
Then for $k\geq 1$ we have
$F\circ\hat{T}^{k-1}(x,\omega)=h(T^{k-1}(x),y_{k})\text{ for }x\in X\text{ and }\omega=(y_{n})_{n\in\mathbb{N}}\in\Omega.$
By the product structure it is straightforward that the sequence of random variables